# ICE-SEARCH: A Language Model-Driven Feature Selection Approach

### v2

@@ 3.2 Evolution-driven Prompting
ICE-SEARCH implements a unique and practical evolution-driven prompting mechanism depicted in Figure 2. This approach allows the LM to iteratively generate and evaluate new feature sets, thereby augmenting its own prompting capabilities. This multi-objective functionality offers several key benefits:
@@
Dynamic Learning and Adaptation: By continually generating and assessing new feature sets while eliminating old feature sets, the backbone LM evolves its understanding of what constitutes an effective feature combination for a given task via in-context learning. This evolutionary learning mechanism ensures that the model remains relevant and creates a continuous improvement loop that augments its prompt over time. Moreover,
Role-played Crossover and Mutation with LMs: A suitably prompted and sampled LM can provide rigorous insights to users’ questions, especially through role-playing the characters of facts-based AI assistants Shanahan et al. (2023). This capability stems from models’ exposure to factual examples and scientifically precise guidances within its training processes Shanahan et al. (2023). Consequently, the agent excels in role-playing as a reliable source of information, drawing on the factual correctness of statistics-driven and fact-based examples in contexts. Hence, ICE Search leverages an LM as the crossover and mutation operator by engaging it with various roles. By integrating LMs directly into the evolutionary search process, we effectively enhance the metrics of interest. This approach allows us to exploit the rich semantic information encoded in LMs, leading to more informed feature selection decisions.
@@
  • Adaptabililty for Diverse Applications: The downstream algorithm, population initialization methods, number of epochs, feature pool size, and the roles assumed by the language model (LM) can all be tailored based on the task’s complexity and specific requirements. Additional parameters, such as early stopping criteria and maximum/minimum number of features, can also be integrated.
@@
  • Scalability to Different Dataset Sizes: Depending on the context length of the backbone LM, our algorithm can be adjusted to handle large-scale datasets that feature thousands or millions of variables, or it can be fine-tuned for smaller datasets to focus on detailed feature analysis.
@@ 3.3 Feature Selection and World Knowledge:
ICE-SEARCH leverages the LM’s extensive world knowledge and cross-domain expertise to facilitate the transfer of insights across related fields. For example, in the context of cardiovascular disease prediction, the algorithm can draw valuable insights from stroke prediction due to its clinical and biological characteristics. Leveraging knowledge from these related fields enhances the feature selection process, potentially revealing common or complementary predictors that hold significance across various health conditions. Using the vast information stored in an LM, ICE-SEARCH looks beyond a dataset and helps end-users identify significant features through analysis of interconnected domains.
@@ 4 Experiments and Evaluation
In this section, we demonstrate ICE-SEARCH’s effectiveness for FS in three tasks: stroke prediction, cardiovascular disease prediction, and diabetes prediction.
We begin by selecting four existing FS methods for initializing the population. The outputs of the four methods then interact with a language model evolutionarily. Note that no finetuning is required in our algorithm, we also examine the algorithm’s generalizability and robustness from experimental results.
@@ 4.1 Public Health Impact and Significance
Fig 3 demonstrates the heightened global and regional impact of HHD, quantified by the Disability-Adjusted Life Year (DALY) per 100,000 persons — one DALY represents the loss of the equivalent of one year of full health (WHO ). The global DALYs attributable to HHD is 2150.8 with a mean of 171.93 and a standard deviation of 267.64 in 2019. They are notably elevated in Asia where DALYs range between 975.46 and 1417.03 (Yang et al. (2023).
The observed geographic disparity in DALYs due to HHD highlights the substantial global health burden it represents, particularly in the hardest-hit regions. This disparity underscores the critical need for effective strategies to identify the causes of HHD for both diagnosis and treatment. Current research is exploring a variety of areas, including risk factors specific to certain demographics, advanced diagnostic techniques, and the genetic basis of these conditions. (NIH ). Hence, any improvement on HHD MPA tasks would save a lot of lives.
@@
Figure 3: Geometric Patterns of DALYs Due to HHD in 2019. (Source of Data: Yang et al. (2023))
@@ 4.3 Experiment Design
In the experiment, we use Phi2 ( Javaheripi and Bubeck (2023)) and 4-bit quantized version of LLaMA2 7B( Touvron et al. (2023)) models respectively as the backbone LM, each containing fewer than 8 billion parameters. The rationale for selecting these LMs lies in their constrained parameter count which allows us to explore the lower boundaries of the performance of ICE-SEARCH.
Initialization: The initialization stage begins by incorporating the backbone LM and four FS algorithms: Decision Tree, Random Forest, Logistic Regression, and Fisher Score Grabczewski and Jankowski (2005); Jaiswal and Samikannu (2017); Cheng et al. (2006); Gu et al. (2012). The LM is subsequently asked to assume the role of a medical doctor to recommend a list of important features based on the given task and dataset features in a zero-shot manner. We prompted the LM five times and the best feature set selected is combined with those selected by the other four FS methods. This combination forms the initial population for our evolution algorithm.
@@
Downstream Model: we employ XGBoost ( Chen and Guestrin (2016)) and Support Vector Machine (SVM) as the downstream models and validated the accuracy of the feature sets through a 10-fold CV to mitigate risks of overfitting. We choose not to conduct an extensive hyperparameter search to emphasize the simplicity and practicality of ICE-SEARCH. The experiments are intended to demonstrate its effectiveness, robustness, generalizability, and potential, rather than claiming it as the state-of-the-art methodology for feature selection.
@@ 4.4 Prompt Design
Zero-Shot Prompting: During initialization, we present scenarios to the model without any prior training on similar tasks via Zero-Shot prompting. For instance, we use prompts like: “Imagine you are a medical doctor. I need you to recommend important features for accurately [TASK DESCRIPTION]. Consider the following features: [FEATURES]. Think step by step. Selected features: …”. The model generate responses by relying solely on its pre-existing knowledge.
Few-Shot Prompting: During the crossover and mutation phases of the algorithm, we adopt a more targeted approach via few-shot prompting. We provide samples of feature sets that had achieved the top five and bottom three validation accuracies in previous assessments. An example prompt for this phase is: “As a Neurologist, recommend important features for [TASK DESCRIPTION]. Consider the following features: [FEATURES]. Here are the selected features and their corresponding classification accuracy results: [DETAILED FEATURE AND ACCURACY LIST]. Be innovative and think step by step. Features selected: …”
Both zero-shot and few-shot prompts, as demonstrated in Fig 2, are crafted to not only direct the model’s attention to specific tasks but also to stimulate explorative and rational thinking.
@@ 4.5.1 Stroke Prediction:
Population Initialization Metric Decision Tree Random Forest Logistic Fisher Score Ensemble of FS ICE-SEARCH
FS methods + Phi2 Test Rank 149.2 48.8 77.4 93.0 51.4 36.4
Val Rank 96.8 8.0 61.6 82.4 2.8 1.0
Worst Test Rank 963.8
Worst Val Rank 977.0
@@
Table 1: Average Ranks (out of 1023) of FS Algorithms on Stroke dataset, using XGBoost as Downstream Classifier and Phi2 as the Backbone LM of ICE-SEARCH.
@@
Table 2: Test Accuracies(%) of FS Algorithms and their ranks (out of 1023) on Stroke dataset, using XGBoost as Downstream classifier and Phi2 as the Backbone LM of ICE-SEARCH. ICES $n^{th}$ stands for the feature set with the $n^{th}$ highest validation accuracy chosen by ICE-SEARCH.
Seed Train Accuracy Validation Accuracy Test Accuracy

### v4

@@ 4.2 Evolution-driven Prompting
ICE-SEARCH implements a unique and practical evolution-driven prompting mechanism depicted in Figure 3. This approach allows the LLM to iteratively generate and evaluate new feature sets, thereby augmenting its own prompting capabilities. This multi-objective functionality offers several key benefits:
@@
Dynamic Learning and Adaptation: By continually generating and assessing new feature sets while eliminating old feature sets, the backbone LLM evolves its understanding of what constitutes an effective feature combination for a given task via in-context learning. This evolutionary learning mechanism ensures that the model remains relevant and creates a continuous improvement loop that augments its prompt over time. Moreover,
Role-played Crossover and Mutation with LLMs: A suitably prompted and sampled LLM can provide rigorous insights to users’ questions, especially through role-playing the characters of facts-based AI assistants( Shanahan et al. (2023)). This capability stems from models’ exposure to factual examples and scientifically precise guidances within its training processes( Shanahan et al. (2023)). Consequently, the agent excels in role-playing as a reliable source of information, drawing on the factual correctness of statistics-driven and fact-based examples in contexts. Hence, ICE Search leverages an LLM as the crossover and mutation operator by engaging it with various roles. By integrating LLMs directly into the evolutionary search process, we effectively enhance the metrics of interest. This approach allows us to exploit the rich semantic information encoded in LLMs, leading to more informed feature selection decisions.
@@
  • Adaptabililty for Diverse Applications: The downstream algorithm, population initialization methods, number of epochs, feature pool size, and the roles assumed by the large language model (LLM) can all be tailored based on the task’s complexity and specific requirements. Additional parameters, such as early stopping criteria and maximum/minimum number of features, can also be integrated.
@@
  • Scalability to Different Dataset Sizes: Depending on the context length of the backbone LLM, our algorithm can be adjusted to handle large-scale datasets that feature thousands or millions of variables, or it can be fine-tuned for smaller datasets to focus on detailed feature analysis.
@@ 4.3 Feature Selection and World Knowledge:
ICE-SEARCH leverages the LLM’s extensive world knowledge and cross-domain expertise to facilitate the transfer of insights across related fields. For example, in the context of cardiovascular disease prediction, the algorithm can draw valuable insights from stroke prediction due to its clinical and biological characteristics. Leveraging knowledge from these related fields enhances the feature selection process, potentially revealing common or complementary predictors that hold significance across various health conditions. Using the vast information stored in an LLM, ICE-SEARCH looks beyond a dataset and helps end-users identify significant features through analysis of interconnected domains.
@@ 5 Experiments and Evaluation
In this section, we demonstrate ICE-SEARCH’s effectiveness for FS in three tasks: stroke prediction ( Fedesoriano (2021)), cardiovascular disease prediction (Ulianova (2019)), and diabetes prediction ( Teboul (2021)).
@@ 5.1 Public Health Impact and Significance
Fig 4 demonstrates the heightened global and regional impact of HHD, quantified by the Disability-Adjusted Life Year (DALY) per 100,000 persons — one DALY represents the loss of the equivalent of one year of full health (WHO (2019)). The global DALYs attributable to HHD is 2150.8 with a mean of 171.93 and a standard deviation of 267.64 in 2019. They are notably elevated in Asia where DALYs range between 975.46 and 1417.03 (Yang et al. (2023b)).
The observed geographic disparity in DALYs due to HHD highlights the substantial global health burden it represents, particularly in the hardest-hit regions. This disparity underscores the critical need for effective strategies to identify the causes of HHD for both diagnosis and treatment. Current research is exploring a variety of areas, including risk factors specific to certain demographics, advanced diagnostic techniques, and the genetic basis of these conditions. (NIH (2021)). Hence, any improvement on HHD MPA tasks would save a lot of lives. As indicated in similar studies, such as Fekadu et al. (2023), high DALYs can be averted as a results of enhanced early treatment likelihood and reduced late treatment-associated mortality. This enhancement helps lower the number of false negatives, thereby reducing the instances of cases that are either missed or diagnosed late. This improvement in diagnosis accuracy can be translated into reduced mortality and DALYs Fekadu et al. (2023). Thus, each percentage improvement in diagnostic accuracy or reduction in errors can translate into significant decreases in DALYs. In fact, if the improved diagnostics lead to a 1% reduction in DALYs, it would equate to 215,000 life years saved globally based on the 2019 DALYs statistics.
@@
Figure 4: Geometric Patterns of DALYs Due to HHD in 2019. (Source of Data: Yang et al. (2023b))
@@ 5.3 Experiment Design
In the experiment, we use Phi2 ( Javaheripi and Bubeck (2023)) and 4-bit quantized version of LLaMA2 7B( Touvron et al. (2023)) models respectively as the backbone LLM, each containing fewer than 8 billion parameters. The rationale for selecting these LLMs lies in their constrained parameter count which allows us to explore the lower boundaries of the performance of ICE-SEARCH.
Initialization: The initialization stage begins by incorporating the backbone LLM and four FS algorithms: Decision Tree, Random Forest, Logistic Regression, and Fisher Score( Grabczewski and Jankowski (2005); Jaiswal and Samikannu (2017); Cheng et al. (2006); Gu et al. (2012)). The LLM is subsequently asked to assume the role of a medical doctor to recommend a list of important features based on the given task and dataset features in a zero-shot manner. We prompted the LLM five times, and the best feature set was combined with those selected by the other four FS methods. This combination forms the initial population for our evolution algorithm.
@@
Downstream Model: we employ XGBoost ( Chen and Guestrin (2016)) and Support Vector Machine (SVM) as the downstream models and validate the accuracy of the feature sets through a 10-fold CV to mitigate risks of overfitting. We choose not to conduct an extensive hyperparameter search to emphasize the simplicity and practicality of ICE-SEARCH. The experiments are intended to demonstrate its effectiveness, robustness, generalizability, and potential rather than claiming it as the state-of-the-art methodology for feature selection.
@@ 5.4 Prompt Design
Zero-Shot Prompting: During initialization, we present scenarios to the model without any prior training on similar tasks via Zero-Shot prompting. For instance, we use prompts like: Imagine you are a medical doctor. I need you to recommend important features for accurately [TASK DESCRIPTION]. Consider the following features: [FEATURES]. Think step by step. Selected features: …”. The model generates responses by relying solely on its pre-existing knowledge.
Few-Shot Prompting: During the crossover and mutation phases of the algorithm, we adopt a more targeted approach via few-shot prompting. We provide samples of feature sets that had achieved the top five and bottom three validation accuracies in previous assessments. An example prompt for this phase is: As a Neurologist, recommend important features for [TASK DESCRIPTION]. Consider the following features: [FEATURES]. Here are the selected features and their corresponding classification accuracy results: [DETAILED FEATURE AND ACCURACY LIST]. Be innovative and think step by step. Features selected: …
Both zero-shot and few-shot prompts, as demonstrated in Fig 3, are crafted to direct the model’s attention to specific tasks and stimulate explorative and rational thinking.
@@ 5.5.1 Stroke Prediction:
Table 1: Average Ranks (out of 1023) of FS Algorithms on Stroke dataset, using XGBoost as Downstream Classifier and Phi2 as the Backbone LLM of ICE-SEARCH.
Metric Decision Tree Random Forest Logistic Fisher Score Ensemble of FS ICE-SEARCH Worst ICE-SEARCH
Test Rank 149.2 48.8 77.4 93.0 51.4 36.4 963.8
Val Rank 96.8 8.0 61.6 82.4 2.8 1.0 977.0
Table 2: Comparative Performance Ranks of ICE-SEARCH on the Stroke Dataset with Different Population Initializations.
Population Initialization Metric ICE-SEARCH
FS methods + Phi2 Test Rank 36.4
Val Rank 1.0
@@
To test this hypothesis, we replace the feature sets chosen by the four traditional feature selection methods with four sets with lower accuracy. The training, validation, and testing accuracies of the four feature sets are provided in Table 13 in Appendix  A.2. As shown in Table 2, these suboptimal feature sets achieve average validation and test accuracy rankings of 38.2 and 1.6, respectively. The results indicate ICE-SEARCH’s ability to target effect feature sets from a baseline of ineffective ones, thus underpinning ICE-SEARCH’s robustness against poor initial feature set selection methods.
Table 3: Test Accuracies(%) of FS Algorithms and their ranks (out of 1023) on Stroke dataset, using XGBoost as Downstream classifier and Phi2 as the Backbone LLM of ICE-SEARCH. ICES $n^{th}$ stands for the feature set with the $n^{th}$ highest validation accuracy chosen by ICE-SEARCH.
@@
Table 4: Average Ranks (out of 1023) of FS Algorithms on Stroke dataset, using Support Vector Machine(SVM) as Downstream Classifier and Phi2 as the Backbone LLM of ICE-SEARCH.
Metric Decision Tree Random Forest Logistic Fisher Score ICE-SEARCH Worst ICE-SEARCH
Test Rank 58.0 7.8 171.2 63.4 13.8 781.2
Val Rank 38.8 7.0 201.2 52.0 1.4 835.8
@@
Table 5: Test Accuracies(%) of FS Algorithms and their ranks (out of 1023) on Stroke dataset, using SVM as Downstream classifier and Phi2 as the Backbone LLM of ICE-SEARCH. ICES $n^{th}$ stands for the feature set with the $n^{th}$ highest validation accuracy chosen by ICE-SEARCH.
Table 6: Average Ranks (out of 1023) of ICE-SEARCH with Different Backbone LLMs on Stroke dataset, using XGBoost as Downstream Classifier.
Metric Phi2 4bit Quantized LLaMA2 7B
Test Rank 36.4 36.4
Val Rank 1.0 1.0
Worst Test Rank 963.88 1022.6
Worst Val Rank 977.0 1023.0
Tables 45, and 6 detail the performance of ICE-SEARCH in stroke prediction, utilizing the Support Vector Machine (SVM) as the downstream classifier and the 4-bit quantized LLaMA 2 7B as the underlying LLM. According to Table 4, although the average test ranking of ICE-SEARCH is 13.8, which is six places lower than the highest average test rank of 7.8, ICE-SEARCH consistently excels in identifying feature sets with the best validation accuracies, achieving an average ranking of 1.4. This performance illustrates ICE-SEARCH’s effectiveness not only in selecting effective feature sets but also in recognizing inadequate ones, as evidenced by the significant deterioration in the worst test and validation ranks from 171.2 and 201.2 to 781.2 and 835.8, respectively—an improvement in identifying less effective feature sets by about 110
To evaluate the ICE-SEARCH algorithm’s adaptability with different underlying Large Language Models (LLMs), its performance was analyzed using a 4-bit quantized LLaMA2 7B. Table 6 offers a comparative analysis of the ICE-SEARCH feature selection (FS) mechanism when applied to Phi2 and LLaMA 2 7B models. Notably, both models achieved identical rankings for test and validation accuracies. However, the 4-bit quantized LLaMA2 7B demonstrated superior capability in identifying less effective feature sets. Specifically, it accurately identified the feature set with the lowest test and validation accuracies, which were ranked at 1022.6 and 1023, respectively.
ICE-SEARCH outperforms conventional feature selection methods in identifying optimal and suboptimal feature sets for stroke prediction. It consistently provides robust and high-quality options across five testing scenarios.
5.5.2 Cardiovascular Disease:
Table 7: Average Ranks (out of 2047) of FS Algorithms on Cardiovascular Disease dataset, using XGBoost as Downstream Classifier and Phi2 as the Backbone LLM of ICE-SEARCH.
Metric Decision Tree Random Forest Logistic Fisher Score Ensemble of Staitstics ICE-SEARCH Worst ICE-SEARCH
Test Rank 34.4 189.8 293.2 53.0 40.0 44.6 1750.2
Val Rank 22.6 202.0 260.4 11.4 4.2 2.8 1761.2
Table 8: Test Accuracies(%) of FS Algorithms and their ranks on Cardiovasclar Disease dataset, using XGBoost as Downstream classifier and Phi2 as the Backbone LLM of ICE-SEARCH. ICES $n^{th}$ stands for the feature set with the $n^{th}$ highest validation accuracy chosen by ICE-SEARCH.
Table 9: Average Test Ranks (out of 2047) of Decision-Randomized ICE-SEARCH on Cardiovascular Disease dataset, using XGBoost as Downstream Classifier and Phi2 as the Backbone LLM of ICE-SEARCH. Seven seeds were chosen as a means to maintain the reproducibility of the randomized decisions. ICES $1^{st}$ stands for the feature set with the highest validation accuracy chosen by ICE-SEARCH.
[…]
96 20.0 56.2
92 40.2 48.4
88 39.4 65.2
84 29.4 31.4
80 29.0 41.6
76 24.0 56.2
The ICE-SEARCH’s performance on Cardiovascular Disease dataset achieves an average rankings of 44.6 for test accuracies and 2.8 for validation accuracies on cardiovascular disease datasets. The Decision-Randomized ICE-SEARCH adaptation enhances feature selection across mismatched datasets and optimizes prediction accuracy with a lower average test accuracy rank in six out of seven cases compared to the standard ICE-SEARCH. This adaptation demonstrates ICE-SEARCH’s adaptability and potential for substantial improvement in prediction scenarios.
5.5.3 Diabetes Prediction:
Table 10: Test Accuracies(%) of FS Algorithms on Diabetes dataset, using XGBoost as Downstream classifier and Phi2 as the Backbone LLM of ICE-SEARCH. ICES $n^{th}$ stands for the feature set with the $n^{th}$ highest validation accuracy chosen by ICE-SEARCH.
42 88.552 90.931 90.613 89.713 90.959 91.002 90.927 90.931 90.897
43 91.017 89.627 90.694 90.752 90.987 91.034 90.992 91.017 90.953
44 90.924 90.87 89.721 90.566 90.93 90.906 90.934 90.87 90.924
45 88.244 90.77 90.715 90.501 90.899 90.77 90.81 90.77 90.747
46 90.949 90.949 90.818 90.099 90.995 90.949 90.818 90.93 90.936
Table 10 presents the comparative performance of FS algorithms on the Diabetes dataset, utilizing XGBoost as a downstream classifier and Phi2 as the LLM. It is important to note that the dataset, after balancing, comprises 436,668 samples and includes 21 indicative variables. This setup enables the examination of ICE-SEARCH’s effectiveness in processing and filtering a large volume of information.
ICE-SEARCH consistently demonstrates superior performance in feature selection and surpasses traditional methods in multiple instances. It consistently achieves higher test accuracies than conventional feature selection methods across various trials, effectively identifying optimal feature sets that provide the most accurate diabetes diagnosis at numerous seeds.
5.6 Rate of Convergence:
ICE-SEARCH algorithm demonstrates fast convergence characteristics in isolating optimal sets of features. A detailed analysis can be found in Appedix A.3
6 Discussion
To effectively improve the metrics of interest and search for valuable feature sets, ICE-SEARCH employs an LLM for crossover and mutation operations and classical FS algorithms for initialization. We have noticed five key advantages of our proposed approach:
High generalizability: ICE-SEARCH can be readily adapted to a variety of tasks and algorithms. As evidenced by Tables (1, 4, 6), ICE-SEARCH continues to output high-performance feature subsets when integrated with different backbone LLMs and downstream predictors, shedding light on the generalizability of ICE-SEARCH. Additionally, Table 10 showscases our algorithm’s ability to manage high-dimensional and high-volume datasets.
Robust against Poor Initialization: As illustrated in Tables (1, 13), ICE-SEARCH has shown robustness to ineffective initialization methods and maintained its feature selection efficiency across three distinct downstream binary classification tasks. It exhibits the capability to adapt to various initialization methods while preserving robustness against the quality of the initial population.
Light-weight training design: ICE-SEARCH leverages the extensive world knowledge pre-trained into an LLM to iteratively refine a population of feature sets. Figs 2 and 3 provide a high-level illustration of our implementtaion of ICE-SEARCH. It employs strategies such as role-playing and few-shot learning without the need for actual fine-tuning. In addition, more advanced techniques such as Tree-of-Thoughts prompting can be applied to the framework ( Yao et al. (2023b)). This training approach effectively reduces the computational demands associated with training high-performance deep learning (DL) models, as highlighted by  Thompson et al. (2020).
7 Conclusion
LLMs with their adeptness at encapsulating vast amounts of world knowledge and facilitating cross-domain insights, have demonstrated exceptional utility across a wide range of fields. This paper introduces ICE-SEARCH, which for the first time explores the integration of LLMs with an evolutionary search framework for FS in MPA tasks.
ICE-SEARCH harnesses the crossover and mutation capabilities inherent in LLMs within an evolutionary framework, significantly improving FS through the model’s comprehensive world knowledge and its adaptability to a variety of roles. Our evaluation of this methodology spans three crucial MPA tasks: stroke, cardiovascular disease, and diabetes, where ICE-SEARCH outperforms traditional FS methods in pinpointing essential features for medical applications.
Appendix A Appendix
A.1 Notations of ICE-Search
Table 11: Summary of Notations used in Algorithm 1
Notation Description
$\mathcal{X}$ A mapping that uses a set of $X$ number of classical FS methods used to initialize the feature set pool. Output $X$ feature sets.
Y Number of times for the LLM to select feature sets through zero-shot prompting for feature set pool initialization.
$\mathcal{Z}$ Set of role-plays for the LLM intended for evolution.
U Number of top feature sets to keep.
V Number of bottom feature sets to keep.
$N$ Number of groups that the given dataset is split into for CV.
$E$ Number of Epochs.
$\mathcal{F}$ The given feature set
$\mathcal{S}$ Feature set pool (i.e. population)
$\mathcal{D}$ Dataset.
$\mathcal{T}$ Task description.
Table 12: Summary of Mapping Functions used in Algorithm 1
Function Description
$\Phi(\mathcal{D},s,N)$ Evaluation of the input feature set $f$ using a pre-determined downstream model on dataset $\mathcal{D}$, output average training and testing accuracies across $N$-fold CV.
[…]
A.2 Poor Population Initialization
Table 13: The accuracy of the four suboptimal feature sets was selected to evaluate the robustness of ICE-SEARCH against poor feature set initialization.
Seed Train Valid Test